OpenStudy (anonymous):

help please f(x)=4x+8

OpenStudy (anonymous):

what the inverse

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

f(x) = 4x-8 y=4x-8 Replace f(x) with y. x=4y-8 Switch y and x and solve for y.

OpenStudy (anonymous):

so y=4x-8

OpenStudy (anonymous):

do i switch x for a random number??

OpenStudy (anonymous):

No, you don't switch it for a random number. Just leave it as it is.

OpenStudy (anonymous):

ok so i got it right

OpenStudy (anonymous):

y=x/4+2

OpenStudy (anonymous):

y=4x-8 4/4x-8 x=2

OpenStudy (anonymous):

oh sry didn't see the response

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

This is what your answer should look like: http://formula.algebra.com/cgi-bin/plot-formula.mpl?expression=f%5Bx%5D%5E-1=%28x%2F4%29%2B2&x=0003

OpenStudy (anonymous):

oh ok thx but i have to write it out step by step

OpenStudy (anonymous):

didn't you already do that f(x) = 4x-8 y=4x-8 Replace f(x) with y. x=4y-8 Switch y and x and solve for y.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

havenice day

OpenStudy (anonymous):

No problem, remember to close this question!